# Emergent predictability in microbial ecosystems

### v2

@@ Acknowledgements.
where $\hat{Y}_{\mu}$ is the model’s prediction of property $Y$ measured in the $\mu$th community and $\beta_{i}^{\Psi}$ are the regression coefficients of each coarse-grained variable (see main text section I.2) to be determined by fitting to a training dataset.
@@
Each regression model corresponding to a coarsened description $\Psi$ is assessed on the same data constructed by randomly splitting datasets into subsets dedicated for either training or testing of regression models. After accounting for replicate communities assayed in each richness tier (i.e., same presence-absence combination $b_{i\mu}$ defined in section I.1), 50% of communities are randomly designated for model training/fitting and the other 50% are reserved for testing/validating prediction error. Model fitting was performed using the standard ordinary least-squares method of MATLAB’s ‘fitlm’ algorithm. The predictions of the fitted model where then compared to the measured values of the test set using standard mean-squared error (MSE). The Fraction of Variance Unexplained (FVU) was computed in the standard way of dividing MSE by the variance of $Y$ across the test set samples. This was repeated for 100 random 50-50 splits of the data, taking the median, upper and lower quartiles across these splits as the values for the ordinate axis and for the error bars around it.
.6 Deducing the Pareto front of efficient coarsened descriptions
The Pareto front of coarsenings $\Psi$ consist of those with minimal prediction error for a given complexity $I(\Psi)$. Ideally, we would like to know this exactly. However, the brute-force approach of performing an exhaustive search through the space of coarse-grainings is computationally infeasible unless the library of strains is very small. This is because the number of possible ways to partition the strains becomes exponentially large: for S strains, the total number is the Bell number $B_{S}$, which is too large to enumerate fully. For example, the 25 taxa used in the Clark et al. already results in $B_{25}\approx 5\times 10^{18}$. This necessitates developing search algorithms to efficiently explore this vast landscape.
With the tractability of a linear regression ansatz, a simple heuristic allows us to construct candidate coarsened descriptions from the microscopic model itself, as we describe below. While these groupings are not guaranteed to be correct or optimal, in practice, we find that they provide good starting points for executing a local search.
Figure S4: A: In the context of a linear regression model class, the coarsening scheme we develop in this work maps onto a regularization scheme. B: This observation suggests a heuristic procedure for constructing candidate coarsenings from the knowledge of the microscopic model (regression with all species kept as separate predictors), by applying a clustering algorithm to the microscopic regression coefficients. In this illustration, there are two clear classes, $\Psi(i)=1$ and $\Psi(i)=2$, from which we can form a candidate two-variable coarsened model with coefficients $\tilde{\beta}_{\alpha}$. Error bars correspond to uncertainty in fit.
Specifically, we note that in the context of linear regression, the coarsening procedure $\Psi$ we consider here (partitioning taxa into groups and summing the abundances of taxa assigned to the same group) can be mapped to a regularization scheme (Figure S4A). This suggests that efficient coarsenings could be constructed by grouping strains with similar regression coefficients (Figure S4B). For this, we implement a generalized K-means clustering algorithm where the cost function accounts for the uncertainty carried by each coefficient estimate: $\mathrm{cost}(\Psi)=\sum_{i}\frac{\beta_{i}-\langle\beta_{j}\rangle_{\alpha_{j}=\alpha_{i}}}{\sigma_{i}}$, where angular brackets denote averaging over all taxa assigned to the same class by the coarsening $\Psi$. This generalized K-means procedure is used to generate candidate groupings for a range of values of $K$, and these groupings are used as seeds for local searches optimizing the Pareto front. The specific steps of the Pareto front search algorithm are as follows:
  • To obtain accurate estimates initially, the microscopic regression models are fit on all the data before deducing the Pareto front coarsenings via this clustering procedure. This provides the initial guess of the Pareto front by the heuristic depicted in Fig. S4.
  • Evaluate the prediction error of all possible single-strain relabelings of strains. That is, for each coarsening $\Psi$ making up the current guess of the Pareto front, generate all neighboring coarsenings that successively re-assign each single strain to all other classes in $\Psi$.
  • Identify any relabeled coarsenings $\Psi$ that result in a median FVU that fall below the convex hall spanned by the lower quartiles in FVU of the current Pareto front coarsenings and update the current guess of the Pareto front by adding these as the new efficient coarsenings.
  • Repeat steps (2) and (3) until no single-strain relabelings of the Pareto front coarsened descriptions fall below the lower FVU quartiles. At this point, the algorithm has converged to at least a local optimum.
.7 No emergent simplicity in predicting total OD as property of interest
In addition to measuring fermentation end products, Clark et al. measured the total optical density (OD) of their communities to compute strain abundances. Applying our framework to the OD measurements as a control, we find that this property does not exhibit any form of emergent simplicity, indicating that the emergent predictability we observe in Fig. 3 is not an artifact of the distributions of strain abundances (see Fig. S5).
Figure S5: Using the total OD measured in the Clark et al. experiments as a property of interest for prediction, we apply our framework to produce the Pareto fronts with description complexity $I(\Psi)$ measured as partition entropy (A) or number of resolved classes (B).
.8 Robustness tests
@@
First, we perform a down-sampling test, shown in Fig. S6. To test robustness to sample size, we down-sampled the low-richness communities to match the number of unique communities sampled in the high-richness dataset. We performed 5 random subsamplings of the low-richness dataset and computed the Pareto fronts for each down-sample. Comparing these in Fig. S6 to the reference high-richness Pareto fronts (purple dashed line), we find that for 4 out of the 5 properties the effect we observe is not due to a smaller amount of data collected on high-richness communities.
@@
Figure S6: Downsampling test: observed emergent predictability is robust to sample size of datasets. To test if emergent predictability is an effect of differing sample sizes, we downsample the low-richness datasets and re-obtain Pareto fronts. If the Pareto fronts inferred from the downsampled datasets do not drop to encompass the higher-richness Pareto fronts, then the predictability that emerges at higher richness is robust to our downsampling test. Each low-richness dataset was randomly sub-sampled to match the number of unique community member combinations assayed in the high-richness dataset. Plots show 5 random sub-samplings in triangles. Low-richness (orange circles) and high-richness (purple dashed line) Pareto fronts from Fig. 3 are shown for comparison. We find that for the majority of properties, the Pareto fronts obtained after downsampling do not close the gap between low richness and high richness, confirming the shift of the Pareto front is a robust observation.
@@
.9 Testing for artifacts
Checking that the emergent predictability observed in Figure 3 is not purely due to chance, we compare the result against Pareto fronts deduced after randomizing the datasets. For designing a stringent randomization test, one seeks to shuffle the data in such a way that breaks the observed effect, while preserving as much of the original structure present in the data. To apply this sentiment to the datasets from [33] and [34], we aim to keep the abundance statistics of each species across communities fixed; so for each strain, we randomly permute its abundance in communities in which the species is initially present. In doing so, we see in Fig. S8 that the emergent predictability in high-richness communities disappears indicated by an increase in prediction error of the high-richness Pareto front obtained for the permuted abundance data. This is similarly broken when we instead permute the measured properties in high-richness communities, shown in Fig. S9. Both tests confirm that the shift of the Pareto front with increasing community richness is not artifact and that the observed emergent predictability is significant.
@@
Figure S8: Permuting abundance data breaks emergent predictability at high richness. To test that the observation of emergent predictability is not due to an artifact in the abundance data, we perform the following permutation test to see if it breaks our observation. We randomize the high-richness communities from the empirical datasets by shuffling the abundances of each strain with other instances in which the strains were present (preserving the presence-absence structure of each community). Pareto fronts are then inferred on the permuted data and plotted using triangle markers. The higher-richness Pareto fronts from Fig. 3 are plotted for reference, as well as the low-richness Pareto fronts for comparison (orange dashed lines). We find that the high-richness Pareto fronts obtained from randomized abundance data do worse than the low-richness Pareto fronts for each property of interest, indicating that our permutation test breaks the observed emergent predictability and confirming the shift of the Pareto fronts is not an artifact.
@@
Figure S9: Permuting property measurements breaks emergent predictability at high richness. To test that the observation of emergent predictability is not due to an artifact in the measurements of properties of interest, we perform the following permutation test to see if it breaks our observation. Similar to Fig. S8, we permuted the property data measured in high-ricnhess communities of the emprical datasets. Pareto fronts are then inferred on the permuted data and plotted using triangle markers. The higher-richness Pareto fronts from Fig. 3 are plotted for reference, as well as the low-richness Pareto fronts for comparison (orange dashed lines). We find that the high-richness Pareto fronts obtained from randomized property measurements do worse than the low-richness Pareto fronts for each property of interest, indicating that our permutation test breaks the observed emergent predictability and confirming the shift of the Pareto fronts is not an artifact.

### v3

@@ B.4 Training and testing of linear regression models
where $\mu$ labels communities, $\hat{Y}_{\mu}$ is the model’s prediction of property $Y$, and $\beta_{i}^{\Psi}$ are the regression coefficients.
To assess the predictive power of a particular coarsened description $\Psi$, the data is randomly split into training and testing subsets. We note that the experimental design of both studies includes replicate communities (i.e., communities with same presence-absence combination $b_{i\mu}$ defined in section I.1). We take care to keep such replicates together, such that 50% of unique compositions are randomly designated for model training, and the other 50% are reserved for evaluating the out-of-sample prediction error.
@@
Model fitting was performed using the standard ordinary least-squares method of MATLAB’s ‘fitlm’ algorithm. We then used the test set to compare model predictions to the measured values. The Fraction of Variance Unexplained (FVU) was computed in the standard way of dividing mean-squared error (MSE) of prediction by the variance of $Y$ across the test set samples. This was repeated for 100 random 50-50 splits of the data; the plots report the median value, and the error bars specify the upper and lower quartiles across these splits.
B.5 Robustness tests
@@
First, we perform a down-sampling test, shown in Fig. S3. To test robustness to sample size, we down-sampled the low-richness communities to match the total number of strain observations across unique community combinations in the high-richness dataset. This ensures each strain is sampled for regression model fitting. We performed 5 random subsamplings of the low-richness dataset and computed the Pareto fronts for each down-sample. We then comparing these curves to the reference high-richness Pareto fronts, to check whether the subsampling might have closed the gap between them. The results are presented in Fig. S3.
For all four observables of Clark et al., subsampling the low-richness data increases stochasticity, but does not significantly shift the Pareto front, confirming that the gap between the high- and low-richness Pareto fronts cannot be attributed to differences in dataset size.
For Kehe et al. (Focal strain abundance), we see that the Pareto fronts computed for the downsampled data vary tremendously between subsampling replicates. To understand this, we note that the extent of subsampling is much more dramatic: the 19433 samples of the low-richness tier are downsampled to only 346 (to match the mid-richness tier); compare to 870 samples (low-richness tier) downsampled to 180 (matching the high-richness tier) for the case of Clark et al. This is compounded with the fact that for Kehe et al., the mid-richness tier is the highest available. As a result, for higher-complexity descriptions the gap either disappears, or is dwarfed by the stochasticity. We note, however, that at low-complexity end, the gap persists (see panel Fig. S3F, which plots the distribution of the FVU values along the transect at 0.35 bits), confirming that emergent predictability is supported even after subsampling.
To formally confirm this, we performed Wilcoxon rank-sum tests between the low- and mid-richness FVU distributions at a given description complexity. Panel Fig. S3G plots the $p$-value of the comparison, and indicates that the high-richness Pareto front remains statistically separable from the ensemble of subsampled low-richness Pareto fronts up to $\sim 2$ bits.
@@
B.6 Testing for artifacts
Checking that the emergent predictability observed in Figure 4 is not purely due to chance, we compare the result against Pareto fronts deduced after randomizing the datasets. For designing a stringent randomization test, one seeks to shuffle the data in such a way that breaks the observed effect, while preserving as much of the original structure present in the data. To apply this to the datasets from [11] and [33], we aim to keep the abundance statistics of each species across communities fixed. For each strain in Ref.[11], we randomly permute its abundance in communities in which the species is initially present. Because Ref.[33] only reports presence-absence of strains, we must instead shuffle by swapping the presence of each strain across a random subset of all possible community samples (i.e., not preserving presence-absence structure) while maintaining the total number of communities in which a given strain is sampled. After doing so, we see that, as expected, the emergent predictability in high-richness communities disappears (Fig. S6), as indicated by an increase in prediction error of the high-richness Pareto front obtained for the permuted abundance data. Emergent predictability is similarly broken when we instead permute the measured properties in high-richness communities, shown in Fig. S7. Both tests confirm that the shift of the Pareto front with increasing community richness is not an artifact of the analysis, and that the observed emergent predictability is significant.
Finally, in addition to measuring fermentation end products, Clark et al. measured the total optical density (OD) of their communities to compute strain abundances. For completeness, Fig. S5 presents the result of applying our framework to the OD measurements. We should emphasize, however, that community OD differs from all other community properties used in this work, since technically, in this case the response variable and the predictor variables are not technically independent. (Recall that for Clark et al, we use OD measurements for scaling the relative abundances profiled by 16S.) In fact, the only reason this analysis is non-trivial is because, as described above, we standardize the abundance data prior to performing the regressions (without this standardization, the total OD would trivially be a sum of input variables, and be perfectly predictable by any coarsened representation). Thus, interpreting Fig. S5 requires extreme caution, and we present it for completeness only. Perhaps the only possible conclusion to draw here is that “not any set of values would be flagged by our framework as exhibiting emergent predictability,” reinforcing the point already demonstrated by the permutation tests above.
B.7 The raw Pareto fronts used in all comparisons
Here, for more direct comparison, we show all the raw Pareto fronts found by our Metropolis search algorithm (see Methods) that were averaged for summary comparisons in the main text. Fig. S8 shows the actual Pareto fronts versus our Gaussian null for quantifying coarse-grainability in Fig. 3. Fig. S9 plots the Pareto fronts for the high- and low-richness communities for models that do not exhibit emergent predictability in Fig. 4. Fig. S10 displays the high- and low-richness Pareto fronts from the physiological feedback model, showing the 10 replicates for 3 feedback strengths that result in varying levels of emergent predictability in Fig. 5C.
Appendix C Other topics
C.1 Structured species pool is not sufficient for emergent predictability
In the main text, we explained that in any real-world setting, (at least some amount of) coarse-grainability is always expected to arise simply because species pool has structure: some species are more similar than others.
However, our proposed argument for the origin of emergent predictability may superficially appear similar (particularly when the model-derived results are rephrased in intuitive general terms in the Discussion). Consider a simple scenario with a highly structured pool of species forming two well-defined classes: A and B. Species of the same class are very similar, but A and B are distinct in important ways. Then, the following word-based argument appears plausible: “As the number of species increases, both classes become represented by many members; thus, within a class, the quirks of precisely which representatives were selected will average out. In contrast, the difference between the two classes will always persist. Thus, as richness increases, one expects the two-group description, A and B, to become increasingly predictive.” This intuition appears to suggest that, just like coarse-grainability, emergent predictability can also arise simply as a consequence of species pool structure, contrary to our observations in the main text.
It is therefore highly instructive to convert this word-based argument into a simple quantitative model. Observing why emergent predictability is not, in fact, automatically observed in this case will not only correct a possible misconception, but will also help clarify the mechanism we propose in Fig. 5.
As the simplest model with such a two-class species pool, we will, once again, adopt a simple ansatz where species are characterized by their contributions $\beta_{i}$ to community property $Y$:
$Y_{\mu}=\sum_{i}\beta_{i}n_{i\mu}\text{, where }\beta_{i}=\begin{dcases}\beta_{A}+\epsilon_{i},&i\in A\\ \beta_{B}+\epsilon_{i},&i\in B.\end{dcases}$
In other words, the contribution is defined primarily by the class identity ($\beta_{A}$ vs. $\beta_{B}$), while $\epsilon_{i}$ (the “quirks”) represent the small deviations of individual strains, and are assumed small. We will take strain $n_{i\mu}$ and $\epsilon_{i}$ to be independent and Gaussian distributed: $n_{i\mu}\sim\mathcal{N}(m_{n},\sigma_{n}^{2})$ and $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, with $\sigma_{\epsilon}^{2}\ll(\beta_{A}-\beta_{B})^{2}$.
The bimodal distribution of coefficients $\beta_{i}$ means that $Y$ will exhibit strong coarse-grainability: the natural two-group description: $n_{A}=\sum_{i\in A}n_{i}$ and $n_{B}=\sum_{i\in B}n_{i}$ has a much lower error than could be expected by chance. We will now consider whether $Y$ also exhibits emergent predictability. For this, we need to compute the fraction of variance of $Y$ explained by the two-group description, and evaluate its dependence on community richness $S=S_{A}+S_{B}$. Here $S_{A}$ and $S_{B}$ denote the the number of strains in the two groups. For simplicity, we will assume that $S_{A}=S_{B}=S/2$, though this can of course be generalized.
Note that the parameter $m_{n}$ can itself depend on $S$ – for example, if community is assembled under a common resource limitation, one may expect $m_{n}\sim 1/S$; similarly, $\sigma_{n}$ may or may not scale with $S$ as well. For our argument below, we will only need to consider their relative scaling. The ratio $\sigma_{n}/m_{n}$ represents the “noisiness” of each individual species abundance. It would be extremely unusual to see this ratio decrease with $S$. Indeed, if some mechanism ensured that the abundance of each individual species becomes increasingly predictable with diversity, this would directly lead to an extremely strong form of emergent predictability, whereby not only the macroscopic community-level properties, but even the microscopic-level description is increasingly predictable. No empirical indications support such a strong form. Thus, in general we expect $\sigma_{n}/m_{n}\sim S^{\alpha}$ with $\alpha\geq 0$. Our argument below will not require any stronger assumptions; this general statement will be sufficient.
Computing the variance of the community property $Y$, we can separate the total variance into two contributions:
$\mathrm{var}_{\mu}[Y_{\mu}]=\underbrace{\mathrm{var}\left[\sum_{i\in A}n_{i\mu}\beta_{A}+\sum_{i\in B}n_{i\mu}\beta_{B}\right]}_{\mathrm{var}_{\mathrm{coarse}}}+\underbrace{\mathrm{var}\left[\sum_{i}n_{i\mu}\epsilon_{i}\right]}_{\mathrm{var}_{\mathrm{quirks}}},$
where the first term is precisely the variance explainable by the coarse-grained description, and the second term is the residual contribution from the individual “quirks.” To test for emergent predictability in $Y$, we need to compute the fraction of the explainable variance relative to the total. Simplifying each term we have
$\mathrm{var}_{\mathrm{coarse}}$ $=(S_{A}\beta_{A}^{2}+S_{B}\beta_{B}^{2})\sigma_{n}^{2}=\frac{1}{2}S\sigma_{n}^{2}\left(\beta_{A}^{2}+\beta_{B}^{2}\right)$
$\mathrm{var}_{\mathrm{quirks}}$ $=(S_{A}+S_{B})\sigma_{\epsilon}^{2}(\sigma_{n}^{2}+m_{n}^{2})=S\sigma_{n}^{2}\sigma_{\epsilon}^{2}(1+(\dots)S^{-2\alpha})\simeq S\sigma_{n}^{2}C\sigma_{\epsilon}^{2}$
In the last line, we used the fact that $\alpha\geq 0$, and therefore the $S^{-2\alpha}$ term in the brackets is either negligible, or at most provides a constant correction of order 1. We conclude that in this example, both contributions to the variance scale in the exact same way with $S$, and the fraction of variance left unexplained by the two-group model remains constant:
$\frac{\mathrm{var}_{\mathrm{quirks}}}{\mathrm{var}_{\mathrm{coarse}}+\mathrm{var}_{\mathrm{quirks}}}=\frac{2C\sigma_{\epsilon}^{2}}{\beta_{A}^{2}+\beta_{B}^{2}+2C\sigma_{\epsilon}^{2}}.$
Therefore, the observable $Y$ exhibits no emergent predictability, despite being strongly coarse-grainable due to species pool structure. Intuitively, the individual strain “quirks” self-average with richness – but so does the variance explained by the coarse model, scaling in the exact same way. Thus, in the absolute sense, the two-class model is an increasingly good one, and this is the phenomenon of self-averaging (unexplained variance decreases). However, the unexplained fraction of variance remains constant (no emergent predictability).
Now, imagine if some mechanism were to maintain a finite range of variability along the A–B axis, such that some communities assembled into mostly A-type strains, while others assembled into mostly B-type strains. Then, as the $\epsilon_{i}$ average out, the variance explained by the class representation would indeed become an increasingly dominant fraction. This is what our Fig. 5 model achieves. In a pH-based intuition: the existence of two classes, acid lovers and acid haters, is not enough. But if there is a feedback that drives community pH in either direction based on initial composition, such that the community alternately assembles into a low-pH state dominated by acid-lovers, or a high-pH state dominated by acid haters – this community will display emergent predictability.
In the actual model we use in Fig. 5, note that turning the $p$-dependent feedback off makes the metabolic activity $a_{i}$ identical for all species at all times; all species “grow as one” until the resource is depleted. Then, it is trivial to see that the contributions to both $p$ and $Y$ experience pure self-averaging, which affects both quantities in the exact same way. In this limit, the failure to observe emergent predictability has the same origin as in the model just described.
C.2 PCA on Clark et al. functional space
The experiments of Ref. [11] measured the final concentration of four fermentation byproducts in each sample community. We can view these measurements as points in a four-dimensional “functional space.” In Fig. 5A, we perform principal component analysis on this functional property data and plot the projections along the first two principal axes for the measurements in low- versus high-richness communities, showing that at higher richness community function is spanned by a lower-dimensional subspace. Fig. S11 quantifies this by plotting the variance explained by each principal component. The plot demonstrates that as richness increases, the first principal axis explains an increasing fraction of total variance, and confirms the numbers quoted in the main text.
Supplementary figures
Figure S1: Redoing the analysis shown in Fig. 4 with description complexity axis $I(\Psi)$ measured in terms of number of classes of a given coarsened description $\Psi$.
Figure S3: Downsampling test: observed emergent predictability is robust to sample size of datasets. To test if our observation of emergent predictability might be affected by the differnces in the number of samples per richness tier, we randomly subsample the low-richness datasets to match the total number of strain observations across the unique high-richness communities and re-obtain Pareto fronts. Plots show 5 random sub-samplings in triangles. Low-richness (orange circles) and high-richness (purple dashed line) Pareto fronts from Fig. 4 are shown for comparison. We find that for the majority of properties, the Pareto fronts obtained after downsampling do not close the gap between low richness and high richness, confirming that the shift of the Pareto front is a robust observation. The result of this test for Focal Strain Abundance (E) appears weaker than the other properties, but computing p-values from Wilcoxon rank-sum tests on the low- and mid-richness FVU distributions across description complexity (G) shows that the two Pareto fronts are statistically separable up to $\sim 2$ bits. Panel F shows an example cross-section at the lowest non-zero description complexity grouping of the mid-richness Pareto, comparing the FVU distributions across cross-validation replicates that constitute the mid-richness and subsampled low-richness fronts. To perform the statistical tests, the description complexity axis was binned at a width of $0.25$ bits and FVU distributions compared within each bin. Red dashed line in G shows the $0.05$ significance threshold.
@@ C.1 Structured species pool is not sufficient for emergent predictability
Figure S5: Using the total OD measured in the Clark et al. experiments as a community property for prediction, we apply our framework to produce the Pareto fronts with description complexity $I(\Psi)$ measured as partition entropy (A) or number of resolved classes (B).
@@
Figure S6: Permuting abundance data breaks emergent predictability at high richness. To test that the observation of emergent predictability is not due to an artifact in the abundance data, we perform the following permutation test to see if it breaks our observation. For Ref. [11] we randomize the high-richness communities by shuffling the abundances of each strain with other instances in which the strains were present (preserving the presence-absence structure of each community). For Ref. [33] we randomize the mid-richness communities by shuffling the presence of each strain while preserving the total number of samples for which a given strain is observed. Pareto fronts are then inferred on the permuted data and plotted using triangle markers; lines with error bars show median and IQR for 5 respective random permutations. The higher-richness Pareto fronts from Fig. 4 are plotted for reference, as well as the low-richness Pareto fronts for comparison (orange dashed lines). We find that the high-richness Pareto fronts obtained from randomized abundance data do worse than the low-richness Pareto fronts for each property of interest, indicating that our permutation test breaks the observed emergent predictability and confirming the shift of the Pareto fronts is not an artifact.
@@
Figure S7: Permuting property measurements breaks emergent predictability at high richness. To test that the observation of emergent predictability is not due to an artifact in the measurements of properties of interest, we perform the following permutation test to see if it breaks our observation. Similar to Fig. S6, we permuted the property data measured in high-ricnhess communities of the emprical datasets. Pareto fronts are then inferred on the permuted data and plotted using triangle markers; lines with error bars show median and IQR for 5 respective random permutations. The higher-richness Pareto fronts from Fig. 4 are plotted for reference, as well as the low-richness Pareto fronts for comparison (orange dashed lines). We find that the high-richness Pareto fronts obtained from randomized property measurements do worse than the low-richness Pareto fronts for each property of interest, indicating that our permutation test breaks the observed emergent predictability and confirming the shift of the Pareto fronts is not an artifact.
Figure S8: Plotting the raw Pareto front data for quantifying coarse-grainability in Fig. 3. Each panel shows the median (line) and IQR (shading) across 100 cross-validation replicates for the actual Pareto versus a Gaussian null (see Methods). Each column corresponds to a different random seed for a given model instance. See text for details on simulation protocols and synthetic community property generation. Row A: The fermentation products measured in Ref. [11] and focal strain abundance measured in Ref. [33]. Rows B,C: Predicting post-invasion abundance of randomly drawn strains in a Consumer-Resource model (CRM) and Generalized Lotka-Volterra (GLV) model. Rows D,E: A synthetic community property produced from a random walk model of an evolved trait on the phylogenetic tree from Ref. [11], with trait conservation depth set to the Genus- (D) or Phylum-level (E). Row F: A synthetic community property generated from applying PICRUSt to KEGG orthologs related to carbon metabolism in the species cultured in Ref. [11]. Fig. 3 is based on averaging across the 10 panels for each model as described in the Methods.
Figure S9: Plotting the raw Pareto front data for quantifying emergent predictability in Fig. 4. Each panel shows the median (line) and IQR (shading) across 100 cross-validation replicates for the high-richness Pareto versus low-richness. Each column corresponds to a different random seed for a given model instance. See text for details on simulation protocols and synthetic community property generation. Rows A,B: Predicting post-invasion abundance of randomly drawn strains in a Consumer-Resource model (CRM) and Generalized Lotka-Volterra (GLV) model. Row C: A synthetic community property produced from a random walk model of an evolved trait on the phylogenetic tree from Ref. [11], with trait conservation depth set to Phylum-level (highly structured case). Row F: A synthetic community property generated from applying PICRUSt to KEGG orthologs related to carbon metabolism in the species cultured in Ref. [11]. Fig. 4 is based on averaging across the 10 panels for each model as described in the Methods.
Figure S10: Plotting the raw Pareto front data for quantifying emergent predictability in the physiological feedback model presented of Fig. 5. Each panel shows the median (line) and IQR (error bars) across 100 cross-validation replicates for the high-richness Pareto versus low-richness. Each column corresponds to a different random seed for a given model instance. See text for details on simulation protocols. Each row corresponds to a different value of feedback strength $\theta$ from the environmental variable pp. Fig. 5 is based on averaging across the 10 panels for each $\theta$ as described in the Methods.
Figure S11: PCA spectra of the four fermentation byproduct concentrations measured in Ref. [11], computed separately in the two tiers of community richness. For high-richness, the dominant component explains $73\%$ of variance, contrasted with $41\%$ for low-richness.
Supplementary tables
Ref. Context $\geq$2 richness tiers? #(samples) / #(strains) Satisfies criteria?
[11] microbial yes 870/25 (low) 477/25 (mid) 180/25 (high) yes
[33] microbial yes 19433/14 (low) 346/14 (mid) yes
[19] microbial yes 21/17 (low) 30/17 (mid) not enough data
[32] grasslands yes 350/31 (low) 200/31 (mid) 50/31 (high) not enough data; not all strains sampled at low and mid richness
[26, 68] grasslands yes 345/32 (low) 22/32 (mid) not enough data at mid richness
[73, 63] grasslands yes 200/60 (low) 40/60 (mid) 10/60 (high) not enough data
[61, 62, 60] grasslands no Not evaluated not enough data; no richness tiers
[70] grasslands yes Not evaluated plots either had all or one species
[30] grasslands no Not evaluated low richness only
Table S1: Catalog of published datasets that assemble communities from a fixed library of species/strains.